### v1

# Understanding the difficulty of low-precision post-training quantization of large language models

Abstract
Large language models of high parameter counts are computationally expensive, yet can be made much more efficient by compressing their weights to very low numerical precision. This can be achieved either through post-training quantization by minimizing local, layer-wise quantization errors, or through quantization-aware fine-tuning by minimizing the global loss function. In this study, we discovered that, under the same data constraint, the former approach nearly always fared worse than the latter, a phenomenon particularly prominent when the numerical precision is very low. We further showed that this difficulty of post-training quantization arose from stark misalignment between optimization of the local and global objective functions. Our findings explains limited utility in minimization of local quantization error and the importance of direct quantization-aware fine-tuning, in the regime of large models at very low precision.
1 Introduction
Large language models (LLMs) have become remarkably powerful in recent years, and are increasingly deployed in real-world applications in practice (Xu et al., 2023; Wang et al., 2024; Zheng et al., 2023; Wu et al., 2023). Despite their effectiveness, present-day LLMs are trained on modern GPUs, with its weights in floating point data type (typically float16). Thus, post-training compression techniques such as quantization, hold promise in making LLMs much more efficient at inference time, crucial for hardware acceleration (Yao et al., 2022; Kim et al., 2023a, b; Park et al., 2022; Frantar et al., 2022a; Lin et al., 2023).
Today, two distinct kinds of methods are popular to quantize an LLM’s weights while maintaining its generalization capabilities. One kind of techniques, namely quantization-aware fine-tuning (QAFT) optimizes a differentiable global objective just like at pretraining time, with quantization operations acting on the model weights; backpropagation and gradient updates are necessary in this case (Chee et al., 2024; Dettmers et al., 2023, 2024; Hu et al., 2021; Li et al., 2023; Huang et al., 2024). The other type of methods seek to minimize layer-wise local errors that are due to the weight quantization; no backpropagation or gradient updates are necessary here (Frantar et al., 2022b; Xiao et al., 2023; Lee et al., 2024; Lin et al., 2023). Note that both approaches require some, typically a small amount of, data to operate. In principle, at the limit where weight quantization error is small, minimizing the local losses should in turn minimize the global loss. Let us first introduce some formal notations before clarification of this point. Denote the LLM network by $f_{{\bm{W}}}:{\mathbb{D}}\to{\mathbb{R}}^{|{\mathcal{V}}|}$, which takes text input ${\bm{x}}\in\mathcal{D}\subset{\mathbb{D}}$ and generates logits $f_{{\bm{W}}}({\bm{x}})$ across vocabulary ${\mathcal{V}}$. Assume it is well-pretrained on language-modeling objective, i.e. ${\bm{W}}=\operatorname*{arg\,min}_{{\bm{W}}^{\prime}}\mathrm{NLL}({\bm{x}}|f_{{\bm{W}}^{\prime}})$ on a training data set $\mathcal{D}$. Here ${\bm{W}}\triangleq\left({\bm{W}}_{1},\cdots,{\bm{W}}_{L}\right)$ collects all layer-wise weights ${\bm{W}}_{l}$ with $l\in\{1,\cdots,L\}$ indexing layers. In the following we will also use vector notation ${\bm{w}}\in{\mathbb{R}}^{D}$ to represent an equivalent, flattened version of ${\bm{W}}$ in the $D$-dimensional weight space. Further denote by $Q:{\mathbb{R}}^{D}\to{\mathbb{R}}^{D}$ the weight (fake-)quantization function (for procedural details see Section 3.2). We refer to the fake-quantized weights as round-to-nearest (RTN), ${\bm{W}}_{\textnormal{RTN}}\triangleq Q({\bm{W}})$.
QAFT methods essentially keep optimizing the global objective with quantization in the loop, i.e.
${\bm{W}}_{\textnormal{QAFT}}$ $=\operatorname*{arg\,min}_{{\bm{W}}^{\prime}}\mathrm{NLL}({\bm{x}}|f_{Q({\bm{W}}^{\prime})}),$ (1)
where ${\bm{x}}\in\mathcal{D}$; due to the non-differentiability of $Q(\cdot)$, straight-through estimator is commonly used in gradient backpropagation.
In contrast, layer-wise quantization error minimization techniques seek to solve $L$ distinct layer-wise optimization problems:
${\bm{W}}_{l}$ $=\operatorname*{arg\,min}_{{\bm{W}}^{\prime}_{l}}\mathrm{MSE}\left(Q({\bm{W}}^{\prime}_{l}){\bm{x}}_{l},{\bm{W}}_{l}{\bm{x}}_{l}\right)$ (2)
$=\operatorname*{arg\,min}_{{\bm{W}}^{\prime}_{l}}\lVert Q({\bm{W}}^{\prime}_{l}){\bm{x}}_{l}-{\bm{W}}_{l}{\bm{x}}_{l}\rVert^{2},$ (3)
where ${\bm{x}}_{l}$ is the input to the $l$-th layer when ${\bm{x}}\in\mathcal{D}$ is passed through the network $f$. One popular method of this kind, namely GPTQ (Frantar et al., 2022a), derived from optimal brain compression (OBC) (Frantar and Alistarh, 2022), seeks to solve the layer-wise mean squared error ($\mathrm{MSE}$) minimization problem by finding a less steeply rising direction in this local quadratic loss landscape, through efficient computation of the Hessian of the local $\mathrm{MSE}$ losses. We denote a GPTQ-optimized network’s weights by ${\bm{W}}_{\textnormal{GPTQ}}$.
Intuitively, minimization of the local $\mathrm{MSE}$ losses as described above should have similar effects as minimizing the global $\mathrm{NLL}$ loss, because at the limit where local $\mathrm{MSE}$s are close to 00, the global $\mathrm{NLL}\left({\bm{x}}\in\mathcal{D_{\mathrm{test}}}|f_{{\bm{W}}_{\textnormal{GPTQ}}}\right)$ should be close to $\mathrm{NLL}\left({\bm{x}}\in\mathcal{D_{\mathrm{test}}}|f_{{\bm{W}}}\right)$, the full-precision pre-trained model. But is this true in practice when local $\mathrm{MSE}$s are not necessarily close to zero?
In this work, by conducting a systematic study on a number of LLMs, we observed and reported a surprising misalignment between the aforementioned two quantization approaches. Specifically, we showed that optimization of the local $\mathrm{MSE}$ losses often could lead to substantially suboptimal global $\mathrm{NLL}$ loss, and a direct minimization of the global $\mathrm{NLL}$ does not necessarily lead to lower local $\mathrm{MSE}$ losses, especially when the quantization precision was very low.
Further, through loss-landscape mapping, we offered an explanation of why the post-training quantizaion by local loss minimization suffered from the difficulty in producing well-generalizing quantized networks at low numerical precision.
We believe our work would inform future LLM quantization practices.
2 Related work
Post-training quantization
Quantization reduces the memory and computational demands of neural networks by converting either their weights or activations from full precision to lower precision formats, like 8-bit integers. Post-training quantization (PTQ) techniques achieve this by representing weights and/or activations with smaller precisions without partially or fully retraining the network. Several PTQ techniques have emerged to reduce deployment cost of LLMs, using different approaches to quantize the models. Some methods focus on identifying outlier features that are difficult to quantize and either represent them with a higher precision, e.g. LLM.int8() (Dettmers et al., 2022), or mitigate their quantization error by adding additional operations to the network, such as SmoothQuant (Xiao et al., 2023), AWQ (Lin et al., 2023), and OWQ (Lee et al., 2024).
Other PTQ methods employ adaptive rounding techniques to compensate for quantization errors. For instance, OBC (Frantar and Alistarh, 2022) quantizes weights one-by-one in a specific order based on the approximate second-order information of the weights, and adjusts the remaining weights to compensate for the quantization error. GPTQ (Frantar et al., 2022a), also known as OPTQ (Frantar et al., 2022b), extends OBC by enabling parallel quantization of weight matrices, applying the same quantization order to all rows of the weight matrix. Similarly, QuIP (Chee et al., 2024) uses an adaptive rounding procedure to minimizes a quadratic proxy objective of the quantization error.
Quantization-aware fine-tuning
Fine-tuning LLMs adapts them to specific tasks, ensuring that they generate accurate and contextually relevant outputs. However, the expensive parameter updates during fine-tuning have drawn significant attention to optimizing its efficiency. Parameter efficient fine-tuning (PEFT) involves reusing some of the pretrained model’s parameters and selectively fine-tuning a subset of parameters for the downstream tasks. Some of PEFT methods are: LoRA (Hu et al., 2021), QLoRA (Dettmers et al., 2024), L4Q (Jeon et al., 2024), LoftQ (Li et al., 2023), Prefix and Prompt Tuning (Li and Liang, 2021; Lester et al., 2021; Qin and Eisner, 2021; Liu et al., 2023), IA3 (Liu et al., 2022), and PEQA (Kim et al., 2024).
LoRA, QLoRA, L4Q, and LoftQ freeze pretrained model parameters and insert task-specific parameters between the frozen layers. These new parameters undergo fine-tuning. Extending this adapter-based approach, the abovementioned methods employ rank decomposition to significantly reduce the number of trainable parameters during fine-tuning (Hu et al., 2021; Dettmers et al., 2024; Jeon et al., 2024; Li et al., 2023). In prefix and prompt tuning, the parameters of an original large language model are frozen, and only the trainable prompt embeddings are fine-tuned (Li and Liang, 2021; Lester et al., 2021; Qin and Eisner, 2021; Liu et al., 2023). Similarly, in IA3 only the hidden state parameters are fine-tuned (Liu et al., 2022).
PEQA is a memory-efficient fine-tuning method for quantized LLMs that updates only the quantization scale, keeping the integer matrix frozen (Kim et al., 2024).
Ultra-low precision pretraining
There has been recent efforts to binarize LLMs by quantizing them to ultra-low bit-widths. For instance, PB-LLM (Shang et al., 2023) and SpQR (Dettmers et al., 2023) employ a mixed-precision quantization technique, representing the majority of the weights with a single bit and retaining a small portion of the weight in the original high precision or int8. BiLLM (Huang et al., 2024) utilize Hessian information to identify salient weights, and non-salients weights. It minimizes the compression loss through an effective binary residual approximation of salient weights and grouped quantization of non-salient weights. BitNet (Wang et al., 2023), on the other hand, replaces the linear layers in the transformer architecture by a binary linear layer with only two weight values, maintaining other components in high-precision. BitNet b1.58 (Ma et al., 2024) is an extension of BitNet that utilizes ternary quantization for its weights. BitNet b1.58 achieves better accuracy in downstream tasks compared to BitNet.
Loss landscape analysis
Analysis of deep neural networks’ loss landscape has long been a tool toward understanding of the generalization properties of the optimized model (Li et al., 2018; Ghorbani et al., 2019; Sagun et al., 2017, 2018), as well as in explaining difficulties arising in efficient network optimization processes (Evci et al., 2020).
3 Methods
3.1 Models and data set
We experimented with $11$ models from $3$ model families, namely GPT-2 (Radford et al., 2019), OPT (Zhang et al., 2022) and Llama 2 (Touvron et al., 2023). All models were served by the Hugging Face Model Hub  [^1] All models were accessed from official repositories hosted by https://huggingface.co/ in April 2024..
We used the WikiText-2 (Merity et al., 2016) data set in all experiments. Unless noted otherwise, the training split used in all procedures was of $128$ examples, each of the maximum sequence length supported by the model being experimented. For test and validation, the entire split was used.
3.2 Numerical data types and quantizer calibration
We experimented with $5$ integer data types for weight quantization, namely int8, int6, int4, int3 and int2, of varied numerical precision. For integer with $B$-bit precision, encoding range is symmetric, i.e. $\{-2^{B-1}+1,\cdots,2^{B-1}-1\}$, excluding $-2^{B-1}$; for example, int2 quantization is effectively ternary, $\{-1,0,1\}$.
We used PyTorch’s quantization API  [^2] See https://pytorch.org/docs/stable/quantization.html. to obtain the weight fake-quantization functions $Q(\cdot)$. The quantization scheme used was per-tensor symmetric  [^3] Note that different quantization schemes and/or different data types of the same precision often result in different generalization quality of the quantized network, determined by the granularity of a channel/group/block-wise scheme. Here we choose to use the simplest scheme in order to conduct a controlled scientific study with fewest confounding factors. . The scaling factor $a_{l}$’s of the fake-quantizer was determined by mean squared quantization error minimization, i.e. $a_{l}=\operatorname*{arg\,min}_{a}\lVert Q_{a}({\bm{W}}_{l})-{\bm{W}}_{l}\rVert^{2},\forall l\in\{1,\cdots\,L\}$, by means of PyTorch’s HistogramObserver mechanism. For simplicity, only weights of layers in the transformer stack were quantized, sparing other weights such as those in embedding and prediction head layers.
Figure 1: A misalignment between minimization of the global $\mathrm{NLL}$ loss (by QAFT) and minimization of the local layer-wise $\mathrm{MSE}$ losses (by GPTQ). Each panel plots results for a single model and for a certain type of loss metric. Panels in the upper row show results of global $\mathrm{NLL}$ loss, whereas those in the lower row of layer-wise local $\mathrm{MSE}$ losses. The three columns correspond to three models whose names are labeled at the top of each panel. All data are plotted for direct comparison between QAFT (vertical axis) and GPTQ (horizontal axis), with the gray diagonal line marking identity. If present, the block dot represents data of the full-precision floating point model. The colored dots chart losses after QAFT and GPTQ, from each of which a colored line emanates and intersects the diagonal, marking the loss of the RTN-quantized model. Each symbol in plots of the lower row represents an individual layer whose weight matrix was subject to quantization.
3.3 GPTQ
We followed the original GPTQ procedure (Frantar et al., 2022a) exactly except for one change described below. As reported by other adopters of GPTQ, e.g. https://huggingface.co/TheBloke, different choices of the dampening factor as a hyperparameter in the GPTQ procedure could lead to outcomes of varied qualities. In order to eliminate this confounding factor, we performed hyperparameter tuning for the GPTQ dampening factor over a search space of $\{10^{-3},\cdots,10^{4}\}$, in a layer-wise manner. As mentioned above, $128$ examples from the training split were used for the GPTQ procedure, consistent with the original work (Frantar et al., 2022a).
3.4 Quantization-aware fine-tuning (QAFT)
To perform QAFT, the exact same $128$ training examples were used as above, for a fair comparison. We ran QAFT for $8$ epochs, i.e. in total $1024$ training iterations (see Appendix B for effect of iterations). Straight through estimator (STE) (Hinton et al., 2012; Bengio et al., 2013) was employed in backpropagation through quantization functions. Only quantized weights in the transformer network were subject to gradient updates. We used the AdamW optimizer (Loshchilov and Hutter, 2017) without weight decay, and with a linear learning rate decay schedule of $1$ order of magnitude. We ran a hyperparameter grid search over $4$ initial learning rate values, $\{10^{-6},10^{-5},10^{-4},10^{-3}\}$ for all models except for Llama-2-7b-hf, in which case it was $\{10^{-6},10^{-3}\}$ (see Appendix A for more details on hyperparameter tuning); best $\mathrm{NLL}$ loss over the validation split was chosen.
4 Results
Figure 2: Tradeoff between quantized model generalization and its weight size. Upper: models from the GPT-2 model family: distilgpt2, gpt2, gpt2-medium, gpt2-large and gpt2-xl. Lower: models from the OPT and Llama 2 families: opt-250m, opt-350m, opt-1.3b, opt-2.7b, opt-6.7b and Llama-2-7b-hf. Black circles represent the full-precision models. Hollow colored circles are RTN-quantized models, solid colored circles QAFT-quantized models, and solid colored squares GPTQ-quantized models. Dotted, dashed and solid gray lines connect quantized solutions from the same model produced by RTN, GPTQ and QAFT, respectively. We highlight the difference between GPTQ- and QAFT-quantized models with colored, transparent, vertical strips, for each quantized model.
Figure 3: Loss landscape analysis of quantized model weights. Data illustrated here are from opt-125m, a network small enough for numerous loss evaluations in order to map the loss landscape in the neighborhood around pretraining convergence. In the legend at the top, we illustrate the mapping strategy in a $2$-dimensional cartoon, which captures key concepts in the $D$-dimensional weight space. The black dot in the middle marks the pretraining convergence ${\bm{w}}$. The continuous loss landscape is probed first by measuring loss at ${\bm{w}}+\lambda\hat{\bm{e}}$, i.e. pretrained weight subject to random perturbation $\hat{\bm{e}}\sim{\mathcal{S}}^{D}$ sampled uniformly from the $D$-dimensional unit sphere. We sweep $\lambda\in{\mathbb{R}}^{+}$ (thin, light gray lines emanating from the black circle) to map the radial loss landscape along a specific random direction $\hat{\bm{e}}$. The gray grid represents the representable weight values prescribed by the weight quantizer $Q(\cdot)$, out of which we show three key quantized weights under question: ${\bm{w}}_{\textnormal{RTN}}=Q({\bm{w}})$ (blue circle), ${\bm{w}}_{\textnormal{QAFT}}$ (green circle), and ${\bm{w}}_{\textnormal{GPTQ}}$ (red circle). We measure the loss function at these key points as well as those along the linear segment resulting from a convex combination of two of these (colored lines). We plot the radial loss landscape ($\mathrm{NLL}$ loss against $\ell_{2}$ distance from ${\bm{w}}$) in the lower panels, training loss on the left and validation loss on the right. Graphical symbols of points and segments are consistent with the legend at the top.
4.1 GPTQ vs. QAFT under the same data constraint
First, we asked whether minimization of the local $\mathrm{MSE}$ losses through GPTQ aligned with minimization of the global $\mathrm{NLL}$ loss through QAFT. In order to answer this question, we conducted the GPTQ and QAFT procedures as described in Section 3 on all experimented LLMs. We reported, for each quantized networks produced by both methods, the global $\mathrm{NLL}$ loss and local layer-wise $\mathrm{MSE}$ losses evaluated on the test data set.
In Figure 1, we plotted the global $\mathrm{NLL}$ and local $\mathrm{MSE}$ losses for gpt2-xl, opt-6.7b and llama2-7b, three example LLMs that were the largest among each model family we experimented with. As evident in the upper row of plots showing global $\mathrm{NLL}$ losses, the colored dots were always below the diagonal identity line, indicating that $\mathrm{NLL}$ loss after QAFT was always lower than that after GPTQ for the same model at a particular precision. At certain precision, sometimes performing GPTQ led to $\mathrm{NLL}$ losses even higher than that of RTN. In contrast, the lower row of plots illustrating layer-wise local $\mathrm{MSE}$ losses showed a remarkably different story. GPTQ always lowered layer-wise $\mathrm{MSE}$ losses, precisely as it was designed to do, but QAFT hardly ever reduced layer-wise $\mathrm{MSE}$s from the RTN level and in most cases increased the layer-wise $\mathrm{MSE}$ losses.
These results suggested that direct minimization of the global $\mathrm{NLL}$ loss did not necessarily lead to a reduction of layer-wise local $\mathrm{MSE}$s. Likewise, indirect minimization of layer-wise $\mathrm{MSE}$s also did not guarantee that the global $\mathrm{NLL}$ loss would decrease, particularly for int2 quantization of most experimented models. Since the generalization capability of a quantized model is measured by the global $\mathrm{NLL}$ loss, minimizing layer-wise $\mathrm{MSE}$s did not seem an ideal surrogate objective function to optimize, in light of the observed misalignment. This effect was especially prominent in the case of low precision formats, where QAFT was remarkably more effective than GPTQ; in Appendix B, we showed that just very few iterations of gradient updates with QAFT could achieve a much better generalizing quantized model than GPTQ.
4.2 Scaling with LLM size and numerical precision
Next, we questioned how the generalization abilities of quantized models produced by RTN, QAFT and GPTQ, scaled as a function of model size and numerical precision. In order to gain a comprehensive understanding of the matter, we aggregated all models in a single tradeoff graph of test $\mathrm{NLL}$ loss plotted against total transformer weight sizes in gigabytes (Figure 2).
As shown in Figure 2, the Pareto front was dominantly occupied by networks quantized through QAFT. In the top panel for the GPT-2 model family, the Pareto front mainly consisted of solutions by QAFT with int6 and int8. At a same model size, a GPT-2 model quantized to int6 and int8 outperformed both a full-precision counterpart with fewer parameters, and one quantized to lower precision with more parameters. We observed a similar pattern with models from the OPT and Llama 2 families, as shown in the lower panel of Figure 2. Here, int4, int6 and int8 models produced by QAFT dominated the Pareto front.
In both panels of Figure 2, we highlighted the difference in test $\mathrm{NLL}$ between QAFT and GPTQ solutions, by the colored, transparent and vertical strips. Evidently, models quantized by QAFT was better generalizing at lower precision, such as int2, int3 and int4. This observation was consistent with the misalignment between the global $\mathrm{NLL}$ loss and the layer-wise local $\mathrm{MSE}$ losses was more prominent at low numerical precision.
4.3 Explanation of the misalignment from a loss landscape perspective
Finally, we asked why minimization of local layer-wise $\mathrm{MSE}$ losses did not align with minimizing the global $\mathrm{NLL}$ loss, especially in case of quantization to low numerical precision. To answer this question, we resorted to mapping of the global $\mathrm{NLL}$ loss landscape in the $D$-dimensional weight space in the neighborhood of the pretrained weights ${\bm{w}}\in{\mathbb{R}}^{D}$.
We presented an example in Figure 3 of opt-125m, a model small enough for numerous loss evaluations.
First, we measured $\mathrm{NLL}$ loss along a number of random directions emanating from ${\bm{w}}$ (the thin, light gray lines in Figure 3). It was evident from the radial mapping that ${\bm{w}}$ sat near the bottom of an attractive basin. The near-convergence neighborhood looked like quadratic, and the loss profiles along randomly sampled directions were rather similar to each other. This is consistent with prior analysis of the loss Hessian spectra showing the presence of a large bulk subspace and a small number of outliers (Sagun et al., 2017, 2018; Ghorbani et al., 2019); thusly, a random linear combination of Hessian eigenvectors would, with high probability, stay within the bulk subspace, making independent radial loss profile look very similar to each other. Beyond the near-convergence locality, notably, the radial loss landscape quickly deviated from a good second-order approximation and plateaued at a high level, demarcating an $\ell_{2}$ neighborhood size $R({\bm{w}})$ that defined the attractive basin.
Next, we charted the loss landscape of the quantized weights via RTN, QAFT and GPTQ, and of linear interpolation segments between them (colored circles and line segments in the Figure 3). We made the following observations.
  • ${\bm{w}}_{\textnormal{RTN}}$ (blue circle), the quantized weight nearest to ${\bm{w}}$, was within the attractive basin for int8 and int6, clearly outside of it for int3 and int2, and around the border for int4.
  • ${\bm{w}}_{\textnormal{GPTQ}}$ (red circle), was significantly more distant from ${\bm{w}}$ than ${\bm{w}}_{\textnormal{RTN}}$, but the direction from ${\bm{w}}$ leading to ${\bm{w}}_{\textnormal{GPTQ}}$ (red segment) was much less steep than any typical directions. The loss values at ${\bm{w}}_{\textnormal{GPTQ}}$ were low for int8 and int6, and high for int4, int3 and int2, not much different from those at ${\bm{w}}_{\textnormal{RTN}}$ for the same precision.
  • ${\bm{w}}_{\textnormal{QAFT}}$ (green circle), was not materially more distant from ${\bm{w}}$ than ${\bm{w}}_{\textnormal{RTN}}$, but at a significantly lower loss level.
  • Notably, all quantized weights more distant from ${\bm{w}}$ than the radius of the attractive basin, did not seem to be simply connected with ${\bm{w}}$. By simply connected, we meant that the loss profile along the linear segment connecting the two is monotonic without a ridge in the middle.
Based on these observations of the analysed loss landscape, we were able to explain the experimental findings reported in previous sections. The minimization of layer-wise local $\mathrm{MSE}$ by GPTQ, as expected, was successful in picking a much less steeply rising direction emanating from ${\bm{w}}$, along which, if the quantized ${\bm{w}}_{\textnormal{GPTQ}}$ was not distant, stayed at a lower loss level, as in the cases for int8 and int6; but if the quantized ${\bm{w}}_{\textnormal{GPTQ}}$ was substantially distant from ${\bm{w}}$, as in the cases for int3 and int2, the initial shallow rise near ${\bm{w}}$ no longer mattered, resulting in very high loss values at the eventual ${\bm{w}}_{\textnormal{GPTQ}}$ instead. In contrast, even at localities distant from ${\bm{w}}$, QAFT still made material progress from ${\bm{w}}_{\textnormal{RTN}}$, leading to ${\bm{w}}_{\textnormal{QAFT}}$ with significantly lower loss, albeit no longer simply connected to ${\bm{w}}$, likely in another distinct attractive basin, reminiscent of similar observations in training of sparse neural networks (Evci et al., 2020).
Our results suggested that, the relative relationship between the sizes of the attractive basin near convergence $R({\bm{w}})$, and of the weight perturbation due to quantization $\lVert\Delta{\bm{w}}\rVert=\lVert{\bm{w}}_{\textnormal{RTN}}-{\bm{w}}\rVert$, seemed to predict how much minimizing local losses and minimizing the global loss would align with each other. When the latter was substantially larger than the former, misalignment invariably occurred.
5 Discussion
In this work, we conducted a systematic study contrasting and comparing the effectiveness of post-training quantization method GPTQ and quantization-aware fine-tuning (QAFT); the former minimized layer-wise local quantization errors whereas the latter minimized the global training loss. We discovered that, with our experimental setting under the same low training data constraint ($128$ training examples), GPTQ was not as effective as QAFT in nearly all cases. Additional experiments showed that, in certain cases, just a few QAFT iterations could overtake GPTQ in producing a better generalizing LLM (see Appendix B).
We found, rather surprisingly, that in the cases where GPTQ fared significantly worse than QAFT, the minimization of the local $\mathrm{MSE}$ and the global $\mathrm{NLL}$ losses did not necessarily align with each other. We further elucidated, through loss landscape analysis, that such misalignment was a characteristic behavior when the weight perturbation due to quantization was significantly larger than the radius of the attractive basin at the pretraining convergence, a frequent situation consequent from very low precision weight quantization.
Our study demonstrated an apparent lack of correlation between the generalization capabilities of a quantized neural network and the local quantization errors in its component layers, the latter of which often used as a metric to evaluate the quality of quantization schemes. We call for caution on generalization of the utility of local error minimization-based post-training quantization techniques in practice, and provided a new perspective toward understanding the difficulty and identification of the circumstances under which applying these methods were effective.
Appendix A Optimal learning rate for QAFT
While performing hyperparameter grid search on initial learning rate over the set $\{10^{-6},10^{-5},10^{-4},10^{-3}\}$, we discovered that the optimal choices strongly correlated with the model size and the quantization precision, which is supported by Fig 4. For a particular model, we observe that QAFT at higher precision such as int8, int6 and int4 favored low learning rates, whereas QAFT at lower precision such as int3 and int2 preferred high learning rates. For a specific format, larger models in the same model family preferred lower learning rates while smaller models in the model family prefers higher learning rates.
Figure 4: Left: validation $\mathrm{NLL}$ loss of gpt2-large after fine-tuning. Dashed lines represent the loss after RTN. Right: best learning rates for GPT-2 family.
Appendix B Effect of number of QAFT iterations
Although we used 8 epochs ($1024$ iterations) to achieve the most optimal performance, we discovered that QAFT for $1$ epoch is sufficient to outperform GPTQ, as shown in Figure 5.
Figure 5: $\mathrm{NLL}$ on test data after each epoch of fine-tuning. Epoch 00 represents the $\mathrm{NLL}$ of RTN, which is the starting point for fine-tuning. The horizontal red line marks the $\mathrm{NLL}$ after GPTQ.
Appendix C Optimal GPTQ dampening factor
In GPTQ, dampening factor is multiplied with the average diagonal value in Hessian matrix $H$ and added to the diagonal entries of $H$ to achieve better numerical stability [Frantar et al., 2022a]. In our experiments, we did a grid search on dampening factor for each layer with the objective of minimizing layer-wise $\mathrm{MSE}$. We visualized the best dampening factor for each quantized layer but failed to find any meaningful patterns.
Appendix D Additional plots for GPTQ vs. QAFT
Appendix E Additional plots for loss landscape
Figure 6: Left: validation loss landscape for opt-350m. Right: validation loss landscape for opt-1.3b. Figure legend consistent with Figure 3.

### v2

@@
-Understanding the difficulty of low-precision post-training quantization of large language models
+Understanding the Difficulty of Low-Precision Post-Training Quantization for LLMs
 Abstract
@@
 We further showed that this difficulty of post-training quantization arose from stark misalignment between optimization of the local and global objective functions.
-Our findings explains limited utility in minimization of local quantization error and the importance of direct quantization-aware fine-tuning, in the regime of large models at very low precision.
-1 Introduction
-Large language models (LLMs) have become remarkably powerful in recent years, and are increasingly deployed in real-world applications in practice (Xu et al., 2023; Wang et al., 2024; Zheng et al., 2023; Wu et al., 2023).
-Despite their effectiveness, present-day LLMs are trained on modern GPUs, with its weights in floating point data type (typically float16).
-Thus, post-training compression techniques such as quantization, hold promise in making LLMs much more efficient at inference time, crucial for hardware acceleration (Yao et al., 2022; Kim et al., 2023a, b; Park et al., 2022; Frantar et al., 2022a; Lin et al., 2023).
-Today, two distinct kinds of methods are popular to quantize an LLM’s weights while maintaining its generalization capabilities.
-One kind of techniques, namely quantization-aware fine-tuning (QAFT) optimizes a differentiable global objective just like at pretraining time, with quantization operations acting on the model weights; backpropagation and gradient updates are necessary in this case (Chee et al., 2024; Dettmers et al., 2023, 2024; Hu et al., 2021; Li et al., 2023; Huang et al., 2024).
-The other type of methods seek to minimize layer-wise local errors that are due to the weight quantization; no backpropagation or gradient updates are necessary here (Frantar et al., 2022b; Xiao et al., 2023; Lee et al., 2024; Lin et al., 2023).
-Note that both approaches require some, typically a small amount of, data to operate.
-In principle, at the limit where weight quantization error is small, minimizing the local losses should in turn minimize the global loss.
-Let us first introduce some formal notations before clarification of this point.
+Our findings suggested limited utility in minimization of local quantization error and the importance of direct quantization-aware fine-tuning, in the regime of large models at very low precision.
+Index Terms:
+Quantization, Large Language Models, Natural Language processing, Efficient Machine Learning
+I Introduction
+Large language models (LLMs) have become remarkably powerful and are increasingly deployed in real-world applications in recent years [42, 39, 45, 40].
+Despite their effectiveness, LLMs are typically trained with weights in float16, making post-training compression techniques such as quantization crucial for efficient inference [43, 17, 18, 31, 9, 24].
+Two distinct methods are prevalent for LLM weight quantization.
+The first one, quantization-aware fine-tuning (QAFT) optimizes a differentiable global objective just like at pretraining time, with quantization operations acting on the model weights, requiring backpropagation and gradient updates [2, 4, 5, 13, 23, 14].
+The second method minimizes layer-wise local quantization errors, with no backpropagation needed [10, 41, 19, 24].
+Both methods typically require a small amount of data and should theoretically achieve similar result since minimizing the local losses should in turn minimize the global loss, and vice versa.
+Let us first introduce some formal notations.
 Denote the LLM network by $f_{{\bm{W}}}:{\mathbb{D}}\to{\mathbb{R}}^{|{\mathcal{V}}|}$, which takes text input ${\bm{x}}\in\mathcal{D}\subset{\mathbb{D}}$ and generates logits $f_{{\bm{W}}}({\bm{x}})$ across vocabulary ${\mathcal{V}}$.
@@
 In the following we will also use vector notation ${\bm{w}}\in{\mathbb{R}}^{D}$ to represent an equivalent, flattened version of ${\bm{W}}$ in the $D$-dimensional weight space.
-Further denote by $Q:{\mathbb{R}}^{D}\to{\mathbb{R}}^{D}$ the weight (fake-)quantization function (for procedural details see Section 3.2).
+We denote $Q:{\mathbb{R}}^{D}\to{\mathbb{R}}^{D}$ as the weight (fake-)quantization function (for procedural details see Section III-B).
 We refer to the fake-quantized weights as round-to-nearest (RTN), ${\bm{W}}_{\textnormal{RTN}}\triangleq Q({\bm{W}})$.
@@
 ${\bm{W}}_{\textnormal{QAFT}}$ $=\operatorname*{arg\,min}_{{\bm{W}}^{\prime}}\mathrm{NLL}({\bm{x}}|f_{Q({\bm{W}}^{\prime})}),$ (1)
-where ${\bm{x}}\in\mathcal{D}$; due to the non-differentiability of $Q(\cdot)$, straight-through estimator is commonly used in gradient backpropagation.
+where ${\bm{x}}\in\mathcal{D}$; due to the non-differentiability of $Q(\cdot)$, straight-through estimator is commonly used in gradient back-propagation.
 In contrast, layer-wise quantization error minimization techniques seek to solve $L$ distinct layer-wise optimization problems:
@@
 where ${\bm{x}}_{l}$ is the input to the $l$-th layer when ${\bm{x}}\in\mathcal{D}$ is passed through the network $f$.
-One popular method of this kind, namely GPTQ (Frantar et al., 2022a), derived from optimal brain compression (OBC) (Frantar and Alistarh, 2022), seeks to solve the layer-wise mean squared error ($\mathrm{MSE}$) minimization problem by finding a less steeply rising direction in this local quadratic loss landscape, through efficient computation of the Hessian of the local $\mathrm{MSE}$ losses.
+One popular method of this kind, namely GPTQ [9], derived from optimal brain compression (OBC) [8], seeks to solve the layer-wise mean squared error ($\mathrm{MSE}$) minimization problem by finding a less steeply rising direction in this local quadratic loss landscape, through efficient computation of the Hessian of the local $\mathrm{MSE}$ losses.
 We denote a GPTQ-optimized network’s weights by ${\bm{W}}_{\textnormal{GPTQ}}$.
-Intuitively, minimization of the local $\mathrm{MSE}$ losses as described above should have similar effects as minimizing the global $\mathrm{NLL}$ loss, because at the limit where local $\mathrm{MSE}$s are close to 00, the global $\mathrm{NLL}\left({\bm{x}}\in\mathcal{D_{\mathrm{test}}}|f_{{\bm{W}}_{\textnormal{GPTQ}}}\right)$ should be close to $\mathrm{NLL}\left({\bm{x}}\in\mathcal{D_{\mathrm{test}}}|f_{{\bm{W}}}\right)$, the full-precision pre-trained model.
-But is this true in practice when local $\mathrm{MSE}$s are not necessarily close to zero?
-In this work, by conducting a systematic study on a number of LLMs, we observed and reported a surprising misalignment between the aforementioned two quantization approaches.
-Specifically, we showed that optimization of the local $\mathrm{MSE}$ losses often could lead to substantially suboptimal global $\mathrm{NLL}$ loss, and a direct minimization of the global $\mathrm{NLL}$ does not necessarily lead to lower local $\mathrm{MSE}$ losses, especially when the quantization precision was very low.
-Further, through loss-landscape mapping, we offered an explanation of why the post-training quantizaion by local loss minimization suffered from the difficulty in producing well-generalizing quantized networks at low numerical precision.
-We believe our work would inform future LLM quantization practices.
-2 Related work
+While intuitively, both approaches should produce well-generalizing quantized networks, our systematic study reveals a misalignment.
+Local loss minimization often leads to suboptimal global loss, and vice versa, especially at low quantization precisions.
+Through loss-landscape analysis, we offered an explanation of why the post-training quantization by local loss minimization struggles to produced well-generalizing quantized networks, guiding future LLM quantization practices.
+II Related work
 Post-training quantization
-Quantization reduces the memory and computational demands of neural networks by converting either their weights or activations from full precision to lower precision formats, like 8-bit integers.
-Post-training quantization (PTQ) techniques achieve this by representing weights and/or activations with smaller precisions without partially or fully retraining the network.
-Several PTQ techniques have emerged to reduce deployment cost of LLMs, using different approaches to quantize the models.
+Quantization reduces the memory and computational demands of neural networks by converting weights or activations from full precision to lower precision formats, like 8-bit integers.
+Post-training quantization (PTQ) techniques achieve this without retraining the network.
+Several PTQ techniques have emerged to reduce deployment cost of LLMs.
 Some methods focus on identifying outlier features that are difficult to quantize and either represent them with a higher precision, e.g.
-LLM.int8() (Dettmers et al., 2022), or mitigate their quantization error by adding additional operations to the network, such as SmoothQuant (Xiao et al., 2023), AWQ (Lin et al., 2023), and OWQ (Lee et al., 2024).
-Other PTQ methods employ adaptive rounding techniques to compensate for quantization errors.
-For instance, OBC (Frantar and Alistarh, 2022) quantizes weights one-by-one in a specific order based on the approximate second-order information of the weights, and adjusts the remaining weights to compensate for the quantization error.
-GPTQ (Frantar et al., 2022a), also known as OPTQ (Frantar et al., 2022b), extends OBC by enabling parallel quantization of weight matrices, applying the same quantization order to all rows of the weight matrix.
-Similarly, QuIP (Chee et al., 2024) uses an adaptive rounding procedure to minimizes a quadratic proxy objective of the quantization error.
+LLM.int8() [3], or mitigate their quantization error by adding additional operations to the network, such as SmoothQuant [41], AWQ [24], and OWQ [19].
+Other PTQ methods employ adaptive rounding techniques to reduce quantization errors.
+For instance, OBC [8] quantizes weights one-by-one in a specific order based on the approximate second-order information of the weights, and adjusts the remaining weights to minimize the quantization error.
+GPTQ [9], also known as OPTQ [10], extends OBC by enabling parallel quantization of weight matrices, applying the same quantization order to all rows of the weight matrix.
+Similarly, QuIP [2] uses adaptive rounding to minimize a quadratic proxy objective of the quantization error.
 Quantization-aware fine-tuning
-Fine-tuning LLMs adapts them to specific tasks, ensuring that they generate accurate and contextually relevant outputs.
-However, the expensive parameter updates during fine-tuning have drawn significant attention to optimizing its efficiency.
-Parameter efficient fine-tuning (PEFT) involves reusing some of the pretrained model’s parameters and selectively fine-tuning a subset of parameters for the downstream tasks.
-Some of PEFT methods are: LoRA (Hu et al., 2021), QLoRA (Dettmers et al., 2024), L4Q (Jeon et al., 2024), LoftQ (Li et al., 2023), Prefix and Prompt Tuning (Li and Liang, 2021; Lester et al., 2021; Qin and Eisner, 2021; Liu et al., 2023), IA3 (Liu et al., 2022), and PEQA (Kim et al., 2024).
-LoRA, QLoRA, L4Q, and LoftQ freeze pretrained model parameters and insert task-specific parameters between the frozen layers.
-These new parameters undergo fine-tuning.
-Extending this adapter-based approach, the abovementioned methods employ rank decomposition to significantly reduce the number of trainable parameters during fine-tuning (Hu et al., 2021; Dettmers et al., 2024; Jeon et al., 2024; Li et al., 2023).
-In prefix and prompt tuning, the parameters of an original large language model are frozen, and only the trainable prompt embeddings are fine-tuned (Li and Liang, 2021; Lester et al., 2021; Qin and Eisner, 2021; Liu et al., 2023).
-Similarly, in IA3 only the hidden state parameters are fine-tuned (Liu et al., 2022).
-PEQA is a memory-efficient fine-tuning method for quantized LLMs that updates only the quantization scale, keeping the integer matrix frozen (Kim et al., 2024).
+Fine-tuning LLMs ensures task-specific adapdations but is computationally expensive.
+Parameter efficient fine-tuning (PEFT) reuses some of the pretrained model’s parameters and selectively fine-tune a subset of parameters for the downstream tasks.
+Common PEFT methods include LoRA [13], QLoRA [5], L4Q [15], LoftQ [23], Prefix and Prompt Tuning [22, 20, 32, 26], IA3 [25], and PEQA [16].
+LoRA, QLoRA, L4Q, and LoftQ freeze pretrained model parameters and fine-tune on inserted task-specific adapters.
+Adapters undergo low rank decomposition to further reduce the trainable parameters [13, 5, 15, 23].
+In prefix and prompt tuning, the parameters of an original large language model are frozen, and only the trainable prompt embeddings are fine-tuned [22, 20, 32, 26].
+Similarly, in IA3 only the hidden state parameters are fine-tuned [25].
+PEQA is a memory-efficient fine-tuning method for quantized LLMs that updates only the quantization scale, keeping the integer matrix frozen [16].
 Ultra-low precision pretraining
-There has been recent efforts to binarize LLMs by quantizing them to ultra-low bit-widths.
-For instance, PB-LLM (Shang et al., 2023) and SpQR (Dettmers et al., 2023) employ a mixed-precision quantization technique, representing the majority of the weights with a single bit and retaining a small portion of the weight in the original high precision or int8.
-BiLLM (Huang et al., 2024) utilize Hessian information to identify salient weights, and non-salients weights.
-It minimizes the compression loss through an effective binary residual approximation of salient weights and grouped quantization of non-salient weights.
-BitNet (Wang et al., 2023), on the other hand, replaces the linear layers in the transformer architecture by a binary linear layer with only two weight values, maintaining other components in high-precision.
-BitNet b1.58 (Ma et al., 2024) is an extension of BitNet that utilizes ternary quantization for its weights.
-BitNet b1.58 achieves better accuracy in downstream tasks compared to BitNet.
+Recent efforts aim to binarize LLMs by quantizing them to ultra-low bit-widths.
+For instance, PB-LLM [36] and SpQR [4] employ a mixed-precision quantization technique, representing the majority of the weights with a single bit while retaining a small portion of the weight in the original high precision or int8.
+BiLLM [14] utilize Hessian information to identify salient and non-sailents weights, employing binary residual approximation of salient weights and grouped quantization of non-salient weights.
+BitNet [38] replaces transformer linear layers by a binary linear layer, retaining other components in high-precision.
+BitNet b1.58 [28] is an extension of BitNet that utilizes ternary quantization for its weights, achieving better accuracy in downstream tasks compared to BitNet.
 Loss landscape analysis
-Analysis of deep neural networks’ loss landscape has long been a tool toward understanding of the generalization properties of the optimized model (Li et al., 2018; Ghorbani et al., 2019; Sagun et al., 2017, 2018), as well as in explaining difficulties arising in efficient network optimization processes (Evci et al., 2020).
-3 Methods
-3.1 Models and data set
-We experimented with $11$ models from $3$ model families, namely GPT-2 (Radford et al., 2019), OPT (Zhang et al., 2022) and Llama 2 (Touvron et al., 2023).
+Analysis of deep neural networks’ loss landscape has long been a tool toward understanding of the generalization properties of the optimized model [21, 11, 34, 35], as well as in explaining difficulties arising in efficient network optimization processes [7].
+III Methods
+III-A Models and data set
+We experimented with $11$ models from $3$ model families, namely GPT-2 [33], OPT [44] and Llama 2 [37].
 All models were served by the Hugging Face Model Hub  [^1] All models were accessed from official repositories hosted by https://huggingface.co/ in April 2024..
-We used the WikiText-2 (Merity et al., 2016) data set in all experiments.
+We used the WikiText-2 [29] data set in all experiments.
 Unless noted otherwise, the training split used in all procedures was of $128$ examples, each of the maximum sequence length supported by the model being experimented.
-For test and validation, the entire split was used.
-3.2 Numerical data types and quantizer calibration
+Entire split was used for validation and test data.
+III-B Numerical data types and quantizer calibration
 We experimented with $5$ integer data types for weight quantization, namely int8, int6, int4, int3 and int2, of varied numerical precision.
@@
 For simplicity, only weights of layers in the transformer stack were quantized, sparing other weights such as those in embedding and prediction head layers.
-Figure 1: A misalignment between minimization of the global $\mathrm{NLL}$ loss (by QAFT) and minimization of the local layer-wise $\mathrm{MSE}$ losses (by GPTQ).
-Each panel plots results for a single model and for a certain type of loss metric.
-Panels in the upper row show results of global $\mathrm{NLL}$ loss, whereas those in the lower row of layer-wise local $\mathrm{MSE}$ losses.
-The three columns correspond to three models whose names are labeled at the top of each panel.
-All data are plotted for direct comparison between QAFT (vertical axis) and GPTQ (horizontal axis), with the gray diagonal line marking identity.
-If present, the block dot represents data of the full-precision floating point model.
-The colored dots chart losses after QAFT and GPTQ, from each of which a colored line emanates and intersects the diagonal, marking the loss of the RTN-quantized model.
-Each symbol in plots of the lower row represents an individual layer whose weight matrix was subject to quantization.
-3.3 GPTQ
-We followed the original GPTQ procedure (Frantar et al., 2022a) exactly except for one change described below.
+Figure 1: Misalignment between minimization of the global $\mathrm{NLL}$ loss (QAFT) and minimization of the local layer-wise $\mathrm{MSE}$ losses (GPTQ).
+The upper row shows global NLL losses, and the lower row presents layer-wise MSE losses for three models (one per column).
+Data points compare QAFT (vertical axis) to GPTQ (horizontal axis).
+The gray diagonal indicates identity.
+Black dots (if present) represent full-precision models, while colored dots mark losses after QAFT and GPTQ.
+Colored lines originating from each dot intersect the diagonal, showing RTN-quantized model losses for the corresponding format.
+In the lower row, symbols represent individual quantized layers.
+III-C GPTQ
+We followed the original GPTQ procedure [9] exactly except for one change described below.
 As reported by other adopters of GPTQ, e.g.
@@
 In order to eliminate this confounding factor, we performed hyperparameter tuning for the GPTQ dampening factor over a search space of $\{10^{-3},\cdots,10^{4}\}$, in a layer-wise manner.
-As mentioned above, $128$ examples from the training split were used for the GPTQ procedure, consistent with the original work (Frantar et al., 2022a).
-3.4 Quantization-aware fine-tuning (QAFT)
+As mentioned above, $128$ examples from the training split were used for the GPTQ procedure, consistent with the original work [9].
+III-D Quantization-aware fine-tuning (QAFT)
 To perform QAFT, the exact same $128$ training examples were used as above, for a fair comparison.
 We ran QAFT for $8$ epochs, i.e.
-in total $1024$ training iterations (see Appendix B for effect of iterations).
-Straight through estimator (STE) (Hinton et al., 2012; Bengio et al., 2013) was employed in backpropagation through quantization functions.
+in total $1024$ training iterations (see Appendix -B for a study on the effect of number of iterations).
+Straight through estimator (STE) [12, 1] was employed in back-propagation through quantization functions.
 Only quantized weights in the transformer network were subject to gradient updates.
-We used the AdamW optimizer (Loshchilov and Hutter, 2017) without weight decay, and with a linear learning rate decay schedule of $1$ order of magnitude.
-We ran a hyperparameter grid search over $4$ initial learning rate values, $\{10^{-6},10^{-5},10^{-4},10^{-3}\}$ for all models except for Llama-2-7b-hf, in which case it was $\{10^{-6},10^{-3}\}$ (see Appendix A for more details on hyperparameter tuning); best $\mathrm{NLL}$ loss over the validation split was chosen.
-4 Results
+We used the AdamW optimizer [27] without weight decay, and with a linear learning rate decay schedule of $1$ order of magnitude.
+We ran a hyperparameter grid search over $4$ initial learning rate values, $\{10^{-6},10^{-5},10^{-4},10^{-3}\}$ for all models except for Llama-2-7b-hf, in which case it was $\{10^{-6},10^{-3}\}$ (see Appendix -A for more details on hyperparameter tuning); best $\mathrm{NLL}$ loss over the validation split was chosen.
+IV Results
 Figure 2: Tradeoff between quantized model generalization and its weight size.
@@
 Black circles represent the full-precision models.
-Hollow colored circles are RTN-quantized models, solid colored circles QAFT-quantized models, and solid colored squares GPTQ-quantized models.
+Hollow colored circles are RTN-quantized models, solid colored circles are QAFT-quantized models, and solid colored squares are GPTQ-quantized models.
 Dotted, dashed and solid gray lines connect quantized solutions from the same model produced by RTN, GPTQ and QAFT, respectively.
@@
 Figure 3: Loss landscape analysis of quantized model weights.
-Data illustrated here are from opt-125m, a network small enough for numerous loss evaluations in order to map the loss landscape in the neighborhood around pretraining convergence.
+Data illustrated here are from opt-125m, a network small enough for numerous loss evaluations.
 In the legend at the top, we illustrate the mapping strategy in a $2$-dimensional cartoon, which captures key concepts in the $D$-dimensional weight space.
@@
 Graphical symbols of points and segments are consistent with the legend at the top.
-4.1 GPTQ vs. QAFT under the same data constraint
-First, we asked whether minimization of the local $\mathrm{MSE}$ losses through GPTQ aligned with minimization of the global $\mathrm{NLL}$ loss through QAFT.
-In order to answer this question, we conducted the GPTQ and QAFT procedures as described in Section 3 on all experimented LLMs.
-We reported, for each quantized networks produced by both methods, the global $\mathrm{NLL}$ loss and local layer-wise $\mathrm{MSE}$ losses evaluated on the test data set.
-In Figure 1, we plotted the global $\mathrm{NLL}$ and local $\mathrm{MSE}$ losses for gpt2-xl, opt-6.7b and llama2-7b, three example LLMs that were the largest among each model family we experimented with.
-As evident in the upper row of plots showing global $\mathrm{NLL}$ losses, the colored dots were always below the diagonal identity line, indicating that $\mathrm{NLL}$ loss after QAFT was always lower than that after GPTQ for the same model at a particular precision.
-At certain precision, sometimes performing GPTQ led to $\mathrm{NLL}$ losses even higher than that of RTN.
-In contrast, the lower row of plots illustrating layer-wise local $\mathrm{MSE}$ losses showed a remarkably different story.
-GPTQ always lowered layer-wise $\mathrm{MSE}$ losses, precisely as it was designed to do, but QAFT hardly ever reduced layer-wise $\mathrm{MSE}$s from the RTN level and in most cases increased the layer-wise $\mathrm{MSE}$ losses.
-These results suggested that direct minimization of the global $\mathrm{NLL}$ loss did not necessarily lead to a reduction of layer-wise local $\mathrm{MSE}$s.
-Likewise, indirect minimization of layer-wise $\mathrm{MSE}$s also did not guarantee that the global $\mathrm{NLL}$ loss would decrease, particularly for int2 quantization of most experimented models.
-Since the generalization capability of a quantized model is measured by the global $\mathrm{NLL}$ loss, minimizing layer-wise $\mathrm{MSE}$s did not seem an ideal surrogate objective function to optimize, in light of the observed misalignment.
-This effect was especially prominent in the case of low precision formats, where QAFT was remarkably more effective than GPTQ; in Appendix B, we showed that just very few iterations of gradient updates with QAFT could achieve a much better generalizing quantized model than GPTQ.
-4.2 Scaling with LLM size and numerical precision
-Next, we questioned how the generalization abilities of quantized models produced by RTN, QAFT and GPTQ, scaled as a function of model size and numerical precision.
-In order to gain a comprehensive understanding of the matter, we aggregated all models in a single tradeoff graph of test $\mathrm{NLL}$ loss plotted against total transformer weight sizes in gigabytes (Figure 2).
-As shown in Figure 2, the Pareto front was dominantly occupied by networks quantized through QAFT.
-In the top panel for the GPT-2 model family, the Pareto front mainly consisted of solutions by QAFT with int6 and int8.
-At a same model size, a GPT-2 model quantized to int6 and int8 outperformed both a full-precision counterpart with fewer parameters, and one quantized to lower precision with more parameters.
-We observed a similar pattern with models from the OPT and Llama 2 families, as shown in the lower panel of Figure 2.
-Here, int4, int6 and int8 models produced by QAFT dominated the Pareto front.
-In both panels of Figure 2, we highlighted the difference in test $\mathrm{NLL}$ between QAFT and GPTQ solutions, by the colored, transparent and vertical strips.
-Evidently, models quantized by QAFT was better generalizing at lower precision, such as int2, int3 and int4.
-This observation was consistent with the misalignment between the global $\mathrm{NLL}$ loss and the layer-wise local $\mathrm{MSE}$ losses was more prominent at low numerical precision.
-4.3 Explanation of the misalignment from a loss landscape perspective
-Finally, we asked why minimization of local layer-wise $\mathrm{MSE}$ losses did not align with minimizing the global $\mathrm{NLL}$ loss, especially in case of quantization to low numerical precision.
-To answer this question, we resorted to mapping of the global $\mathrm{NLL}$ loss landscape in the $D$-dimensional weight space in the neighborhood of the pretrained weights ${\bm{w}}\in{\mathbb{R}}^{D}$.
-We presented an example in Figure 3 of opt-125m, a model small enough for numerous loss evaluations.
-First, we measured $\mathrm{NLL}$ loss along a number of random directions emanating from ${\bm{w}}$ (the thin, light gray lines in Figure 3).
-It was evident from the radial mapping that ${\bm{w}}$ sat near the bottom of an attractive basin.
-The near-convergence neighborhood looked like quadratic, and the loss profiles along randomly sampled directions were rather similar to each other.
-This is consistent with prior analysis of the loss Hessian spectra showing the presence of a large bulk subspace and a small number of outliers (Sagun et al., 2017, 2018; Ghorbani et al., 2019); thusly, a random linear combination of Hessian eigenvectors would, with high probability, stay within the bulk subspace, making independent radial loss profile look very similar to each other.
-Beyond the near-convergence locality, notably, the radial loss landscape quickly deviated from a good second-order approximation and plateaued at a high level, demarcating an $\ell_{2}$ neighborhood size $R({\bm{w}})$ that defined the attractive basin.
-Next, we charted the loss landscape of the quantized weights via RTN, QAFT and GPTQ, and of linear interpolation segments between them (colored circles and line segments in the Figure 3).
-We made the following observations.
-  • ${\bm{w}}_{\textnormal{RTN}}$ (blue circle), the quantized weight nearest to ${\bm{w}}$, was within the attractive basin for int8 and int6, clearly outside of it for int3 and int2, and around the border for int4.
-  • ${\bm{w}}_{\textnormal{GPTQ}}$ (red circle), was significantly more distant from ${\bm{w}}$ than ${\bm{w}}_{\textnormal{RTN}}$, but the direction from ${\bm{w}}$ leading to ${\bm{w}}_{\textnormal{GPTQ}}$ (red segment) was much less steep than any typical directions.
-The loss values at ${\bm{w}}_{\textnormal{GPTQ}}$ were low for int8 and int6, and high for int4, int3 and int2, not much different from those at ${\bm{w}}_{\textnormal{RTN}}$ for the same precision.
-  • ${\bm{w}}_{\textnormal{QAFT}}$ (green circle), was not materially more distant from ${\bm{w}}$ than ${\bm{w}}_{\textnormal{RTN}}$, but at a significantly lower loss level.
-  • Notably, all quantized weights more distant from ${\bm{w}}$ than the radius of the attractive basin, did not seem to be simply connected with ${\bm{w}}$.
-By simply connected, we meant that the loss profile along the linear segment connecting the two is monotonic without a ridge in the middle.
-Based on these observations of the analysed loss landscape, we were able to explain the experimental findings reported in previous sections.
-The minimization of layer-wise local $\mathrm{MSE}$ by GPTQ, as expected, was successful in picking a much less steeply rising direction emanating from ${\bm{w}}$, along which, if the quantized ${\bm{w}}_{\textnormal{GPTQ}}$ was not distant, stayed at a lower loss level, as in the cases for int8 and int6; but if the quantized ${\bm{w}}_{\textnormal{GPTQ}}$ was substantially distant from ${\bm{w}}$, as in the cases for int3 and int2, the initial shallow rise near ${\bm{w}}$ no longer mattered, resulting in very high loss values at the eventual ${\bm{w}}_{\textnormal{GPTQ}}$ instead.
-In contrast, even at localities distant from ${\bm{w}}$, QAFT still made material progress from ${\bm{w}}_{\textnormal{RTN}}$, leading to ${\bm{w}}_{\textnormal{QAFT}}$ with significantly lower loss, albeit no longer simply connected to ${\bm{w}}$, likely in another distinct attractive basin, reminiscent of similar observations in training of sparse neural networks (Evci et al., 2020).
-Our results suggested that, the relative relationship between the sizes of the attractive basin near convergence $R({\bm{w}})$, and of the weight perturbation due to quantization $\lVert\Delta{\bm{w}}\rVert=\lVert{\bm{w}}_{\textnormal{RTN}}-{\bm{w}}\rVert$, seemed to predict how much minimizing local losses and minimizing the global loss would align with each other.
-When the latter was substantially larger than the former, misalignment invariably occurred.
-5 Discussion
-In this work, we conducted a systematic study contrasting and comparing the effectiveness of post-training quantization method GPTQ and quantization-aware fine-tuning (QAFT); the former minimized layer-wise local quantization errors whereas the latter minimized the global training loss.
-We discovered that, with our experimental setting under the same low training data constraint ($128$ training examples), GPTQ was not as effective as QAFT in nearly all cases.
-Additional experiments showed that, in certain cases, just a few QAFT iterations could overtake GPTQ in producing a better generalizing LLM (see Appendix B).
-We found, rather surprisingly, that in the cases where GPTQ fared significantly worse than QAFT, the minimization of the local $\mathrm{MSE}$ and the global $\mathrm{NLL}$ losses did not necessarily align with each other.
-We further elucidated, through loss landscape analysis, that such misalignment was a characteristic behavior when the weight perturbation due to quantization was significantly larger than the radius of the attractive basin at the pretraining convergence, a frequent situation consequent from very low precision weight quantization.
-Our study demonstrated an apparent lack of correlation between the generalization capabilities of a quantized neural network and the local quantization errors in its component layers, the latter of which often used as a metric to evaluate the quality of quantization schemes.
-We call for caution on generalization of the utility of local error minimization-based post-training quantization techniques in practice, and provided a new perspective toward understanding the difficulty and identification of the circumstances under which applying these methods were effective.
-Appendix A Optimal learning rate for QAFT
-While performing hyperparameter grid search on initial learning rate over the set $\{10^{-6},10^{-5},10^{-4},10^{-3}\}$, we discovered that the optimal choices strongly correlated with the model size and the quantization precision, which is supported by Fig 4.
-For a particular model, we observe that QAFT at higher precision such as int8, int6 and int4 favored low learning rates, whereas QAFT at lower precision such as int3 and int2 preferred high learning rates.
-For a specific format, larger models in the same model family preferred lower learning rates while smaller models in the model family prefers higher learning rates.
-Figure 4: Left: validation $\mathrm{NLL}$ loss of gpt2-large after fine-tuning. Dashed lines represent the loss after RTN. Right: best learning rates for GPT-2 family.
-Appendix B Effect of number of QAFT iterations
-Although we used 8 epochs ($1024$ iterations) to achieve the most optimal performance, we discovered that QAFT for $1$ epoch is sufficient to outperform GPTQ, as shown in Figure 5.
-Figure 5: $\mathrm{NLL}$ on test data after each epoch of fine-tuning.
-Epoch 00 represents the $\mathrm{NLL}$ of RTN, which is the starting point for fine-tuning.
-The horizontal red line marks the $\mathrm{NLL}$ after GPTQ.
-Appendix C Optimal GPTQ dampening factor
-In GPTQ, dampening factor is multiplied with the average diagonal value in Hessian matrix $H$ and added to the diagonal entries of $H$ to achieve better numerical stability [Frantar et al., 2022a].
-In our experiments, we did a grid search on dampening factor for each layer with the objective of minimizing layer-wise $\mathrm{MSE}$.
-We visualized the best dampening factor for each quantized layer but failed to find any meaningful patterns.
-Appendix D Additional plots for GPTQ vs. QAFT
-Appendix E Additional plots for loss landscape
-Figure 6: Left: validation loss landscape for opt-350m. Right: validation loss landscape for opt-1.3b. Figure legend consistent with Figure 3.
+IV-A GPTQ vs. QAFT under the same data constraint
+First, we investigated whether minimization of the local $\mathrm{MSE}$ losses through GPTQ aligned with minimization of the global $\mathrm{NLL}$ loss through QAFT.
+Both methods were applied on all experimented LLMs, as described in Section III, and the resulting global $\mathrm{NLL}$ and local layer-wise $\mathrm{MSE}$ losses were evaluated on the test dataset.
+Figure 1 compares the global $\mathrm{NLL}$ and local $\mathrm{MSE}$ losses for gpt2-xl, opt-6.7b and llama2-7b, which were the largest models among each model family we experimented with.
+In the upper rows, global $\mathrm{NLL}$ loss for QAFT are consistently lower than those for GPTQ, as indicated by the position of the colored dots below the diagonal identity line.
+In some cases, GPTQ even resulted in higher $\mathrm{NLL}$ losses than RTN.
+In contrast, the lower row shows that GPTQ always reduced layer-wise $\mathrm{MSE}$ losses as designed, whereas QAFT maintains or even increases the $\mathrm{MSE}$ losses from RTN.
+These results indicate a misalignment: minimizing global $\mathrm{NLL}$ loss via QAFT does not necessarily reduce local $\mathrm{MSE}$, and minimizing local $\mathrm{MSE}$ loss via GPTQ does not necessarily reduce global $\mathrm{NLL}$.
+This misalignment is particularly evident in low-precision formats, where QAFT significantly outperforms GPTQ.
+Since the generalization capability of a quantized model is measured by the global $\mathrm{NLL}$ loss, minimizing layer-wise $\mathrm{MSE}$s seems to be an ineffective surrogate, in light of the observed misalignment.
+In Appendix -B, we showed that even a few QAFT iterations can produce better generalizing quantized model than GPTQ.
+IV-B Scaling with LLM size and numerical precision
+Next, we examined how the generalization abilities of quantized models produced by RTN, QAFT and GPTQ, scaled with model size and numerical precision.
+To capture this, we plotted test $\mathrm{NLL}$ loss against total transformer weight size (in gigabytes) for all models on a single trade-off graph (Figure 2).
+As shown in Figure 2, QAFT consistently dominated the Pareto front across model families.
+For GPT-2 models, QAFT quantized solutions at int6 and int8 occupied the Pareto front, outperforming both smaller full-precision models and larger models quantized with lower precisions.
+Similar trends were observed for the OPT and Llama 2 families, where QAFT quantized models at int4, int6, and int8 dominated.
+Colored vertical strips in Figure 2 highlight test $\mathrm{NLL}$ differences between QAFT and GPTQ.
+Models quantized by QAFT outperforms GPTQ more significantly at lower precisions, such as int2, int3 and int4, aligning with the earlier observation that misalignment between global $\mathrm{NLL}$ and local $\mathrm{MSE}$ losses is more pronounced at low numerical precision.
+IV-C Explanation of the misalignment from a loss landscape perspective
+Finally, we investigated why minimization of local layer-wise $\mathrm{MSE}$ losses did not align with minimizing the global $\mathrm{NLL}$ loss, especially at low numerical precision.
+To address this, we analyzed the global $\mathrm{NLL}$ loss landscape in the $D$-dimensional weight space around the pretrained weights ${\bm{w}}\in{\mathbb{R}}^{D}$.
+We measured $\mathrm{NLL}$ loss along a number of random directions emanating from ${\bm{w}}$ (the thin, light gray lines in Figure 3).
+The radial mapping revealed that ${\bm{w}}$ sat near the bottom of an attractive basin.
+Within this near-convergence region, the loss landscape appeared quadratic, with similar profiles across various random directions.
+This aligns with prior analysis of the loss Hessian spectra showing a dominant bulk subspace with a few outliers  [34, 35, 11].
+Hence, a random linear combination of Hessian eigenvectors would likely stay within the bulk subspace, resulting in similar radial profiles.
+Beyond the near-convergence locality, the loss landscape deviated from a quadratic approximation and plateaued at a high level, defining the attractive basin’s radius $R({\bm{w}})$.
+Next, we charted the loss landscape of the quantized weights via RTN, QAFT and GPTQ, and of linear interpolation segments between them (colored circles and line segments in Figure 3).
+Key observations include:
+  • ${\bm{w}}_{\textnormal{RTN}}$ (blue circle): The closest quantized weight to ${\bm{w}}$, It was within the attractive basin for int8 and int6, near the border for int4, and outside it for int3 and int2.
+  • ${\bm{w}}_{\textnormal{GPTQ}}$ (red circle): Further from ${\bm{w}}$ than ${\bm{w}}_{\textnormal{RTN}}$, but along a flatter loss direction (red segment).
+The loss values at ${\bm{w}}_{\textnormal{GPTQ}}$ were low for int8 and int6, but high for int4, int3 and int2, similar to ${\bm{w}}_{\textnormal{RTN}}$ at the same precision.
+  • ${\bm{w}}_{\textnormal{QAFT}}$ (green circle): Slightly further from ${\bm{w}}$ than ${\bm{w}}_{\textnormal{RTN}}$, yet achieving significantly lower loss level.
+  • Connectivity: Quantized weights beyond the basin’s radius were not simply connected to ${\bm{w}}$.
+Linear interpolation between ${\bm{w}}$ and these weights showed non-monotonic loss profiles with ridges in the middle.
+Based on these observations, we are able to explain experimental findings from previous sections.
+Minimizing layer-wise local $\mathrm{MSE}$ by GPTQ effectively identified less steeply rising directions from ${\bm{w}}$, resulting in lower loss levels when ${\bm{w}}_{\textnormal{GPTQ}}$ remained within the attractive basin (e.g.
+int8 and int6).
+However, for larger quantization-induced perturbations, the shallow rise near ${\bm{w}}$ is insufficient, resulting in high loss values at ${\bm{w}}_{\textnormal{GPTQ}}$ (e.g.
+int3 and int2).
+In contrast, QAFT consistently follow less steep directions from ${\bm{w}}_{\textnormal{RTN}}$, even when ${\bm{w}}_{\textnormal{QAFT}}$ was distance from ${\bm{w}}$.
+This led to significantly lower loss levels, albeit in a separate attractive basin, similar to patterns observed in sparse networks [7].
+Our results suggested that the alignment between local and global loss minimization depends on the relationship between the attractive basin size $R({\bm{w}})$ and the quantization-induced weight perturbation $\lVert\Delta{\bm{w}}\rVert=\lVert{\bm{w}}_{\textnormal{RTN}}-{\bm{w}}\rVert$.
+When $\lVert\Delta{\bm{w}}\rVert$ is substantially greater than $R({\bm{w}})$, misalignment occurs.
+IV-D Generalization on different datasets
+To assess the generalizability of our findings, we repeated a subset of experiments on C4 [6] and LAMBADA [30] datasets.
+In both cases, we observed consistent misalignment between the minimization of global and local loss (Figure 4), as well as similar trends in the loss landscape (Figure 5).
+Figure 4: Misalignment between minimization of the global NLL loss (by QAFT) and minimization of the local layer-wise MSE losses (by GPTQ) on different datasets.
+Follows the same conventions as Figure 1.
+Figure 5: Loss landscape analysis of quantized model weights on different datasets. Plots showing results for opt-125m quantized in int4. Follows the same conventions as Figure 3.
+V Discussion
+In this work, we systematically compared the effectiveness of post-training quantization GPTQ and quantization-aware fine-tuning (QAFT), where GPTQ minimized layer-wise local quantization errors and QAFT minimized the global training loss.
+Under the same low training data constraint ($128$ training examples), QAFT consistently outperformed GPTQ.
+Surprisingly, we observed that GPTQ’s local $\mathrm{MSE}$ minimization and QAFT’s global $\mathrm{NLL}$ minimization often misalign, especially pronounced at very low precision quantization.
+Through loss landscape analysis, we elucidated that such misalignment was prominent when the weight perturbation due to quantization was significantly larger than the radius of the attractive basin at the pretraining convergence.
+Our findings reveal a lack of correlation between a quantized network’s generalization ability and its local quantization errors, challenging the common reliance on local error minimization metrics for evaluating quantization schemes.
+We urge caution in generalizing the utility of local-error-based post-training quantization and provided a new perspective for understanding the difficulty and identification of conditions where these methods are effective.